Santorino Company produces two models of a component, Model K-3 and Model P-4. The unit contribution margin for Model K-3 is $6; the unit contribution margin for Model P-4 is $14. Each model must spend time on a special machine. The firm owns two machines that together provide 4,000 hours of machine time per year. Model K-3 requires 15 minutes of machine time; Model P-4 requires 30 minutes of machine time.

What is the amount of machine time for model P-4 in terms of percent of a machine hour?

a.

50%

b.

10%

c.

25%

d.

20%

e.

30%
Answer #1

Machine time for Model P-4 = 30 minutes = 0.50 hour

In percentage, Machine time = 0.50/1 = 50% of a machine hour